OpenStudy (anonymous):

An urn contains nine red balls, nine white balls, and nine blue balls, and sample of four balls is drawn at random without replacement. Compute the probability that all of the balls in the sample are the same color.

OpenStudy (anonymous):

so C(27,4) would give the formula for the total samples. Which leads to the formula (27!)/4!(27-4)! = 17550 samples. From here I find the possible value of choosing all colors of the same type. I start by finding the possible ways of having a single color. C(9,4)= (9!)/4!(9-4)!=126 So there are 126 combinations, lets say of red. Since we have 3 colors, i multiple 126 by 3 = 378/17550 = .0215